As CFO of Bond Corporation you have some profits that you wish to hold as retained earning for the next couple of years (and will then be used to fund some major projects that are in development stage). You have to advise the Board on how to invest these earnings and you are considering possible share purchases around which you have collected the following market intelligence:

・Shares of Chyrox Inc recently paid a dividend of $1.50 and dividends are expected to grow at an annual rate of 20 percent for the next 5 years.  A market advisor suggests that the shares have a required rate of return of 15 percent and expects the company to grow at 20 percent per annum for the for-seeable future (though you expect it to come closer to the industry growth rate of around 5% after 5 years).

・The earnings of Foggy Futures Weather Forecasting Company are expected to grow at an annual rate of 14% over the next 5 years and then slow to a constant rate of 6% per year. Foggy currently pays a dividend of $0.36 per share. The CAPM model suggests that the investment needs a 16% rate of return?

Based on this market research what price would you advise your client to be the maximum at which they should buy? Also briefly explain why the market advisor’s expectation beyond 5 years for Chyrox is not realistic.
Answer #1
SHARES OF CHYROX INC:
Year Dividend PVIF ar 15% PV at 15%
0 $             1.50 1
1 $             1.80 0.86957 $         1.57
2 $             2.16 0.75614 $         1.63
3 $             2.59 0.65752 $         1.70
4 $             3.11 0.57175 $         1.78
5 $             3.73 0.49718 $         1.86
PV of dividends t1 to t5 $         8.54
Terminal value of dividends at t5 = 3.73*1.05/(0.15-0.05) = $    39.17
PV of terminal value = 39.17*0.49718 = $       19.47
Price of the share = 8.54+19.47 = $       28.01
SHARES OF FOGGY FUTURES:
Year Dividend PVIF ar 16% PV at 16%
0 $             0.36 1
1 $             0.41 0.86207 $         0.35
2 $             0.47 0.74316 $         0.35
3 $             0.53 0.64066 $         0.34
4 $             0.61 0.55229 $         0.34
5 $             0.69 0.47611 $         0.33
PV of dividends t1 to t5 $         1.71
Terminal value of dividends at t5 = 0.69*1.06/(0.16-0.06) = $       7.31
PV of terminal value = 7.31*0.47611 = $         3.48
Price of the share = 1.71+3.48 = $         5.19
ANSWER:
The maximum price that can be paid is $28.01 for a share of
Chyrox Inc and $5.19 for a share of Foggy Futures.
The expectation of the market advisor as to the growth rate
of Chyrox beyond the 5th year is not realistic, as the industry's growth is expected
to slow down to 6% after the fifth year. It would generally be
impossible for a firm to have higher growth rate once the industry's growth rate
slows down to a constant %.